Determine the length of the arc (in radian measure) and the measure of the angle (in radian degree measures) generated by a point that starts (1,0)and terminates at the following:




1. Positive x-axis



2. Negative x-axis



3.Positive y-axis



4.Negative y-axis

1
Expert's answer
2021-12-13T17:41:58-0500

The circumstance "(2\u03c0r)" of a unit circle is "2\u03c0(1) = 2\u03c0."

1. An angle starting at Cartesian point "(1,0)" and terminating at the positive "x" - axis would represent angle of "2\pi" radians, this would be complete circumstance arc length, that is "2\pi" radians. We can consider the angles of "2\pi n" radian, "n\in \N" and the corresponding lengths of arc of "2\pi n" unis of length.


2. An angle starting at Cartesian point "(1,0)" and terminating at the negative "x" - axis would represent angle of "\pi" radians, this would be complete the half of circumstance arc length, that is "\pi" radians. We can consider the angles of "\pi+2\pi n" radian, "n\in \N" and the corresponding lengths of arc of "\pi+2\pi n" unis of length.


3. An angle starting at Cartesian point "(1,0)" and terminating at the positive "y" - axis would represent angle of "\pi/2" radians, this would be complete the 1/4 of circumstance arc length, that is "\pi/2" radians. We can consider the angles of "\pi/2+2\pi n" radian, "n\in \N" and the corresponding lengths of arc of "\pi/2+2\pi n" unis of length.


4. An angle starting at Cartesian point "(1,0)" and terminating at the negative "y" - axis would represent angle of "3\pi/2" radians, this would be complete the 3/4 of circumstance arc length, that is "3\pi/2" radians. We can consider the angles of "3\pi/2+2\pi n" radian, "n\in \N" and the corresponding lengths of arc of "3\pi/2+2\pi n" unis of length.